The present capital structure is as follows;


2 000 000 ordinary shares with a par value of R1.00 per share. These shares are


currently trading at R2.50 per share and the latest dividend paid is 40 cents. An


average dividend growth of 9% is maintained.


1 500 000 8% R2.00 preference shares, with a market value of R1.80 per share.R10 000 000 non- distributable reserves


R2 000 000 7% debentures due in 6 years time and the current yield to maturity is


10%, and R1 000 000 15% bank loan.


Additional information:


The company has a beta of 2.1 a risk free rate of 7% and a return of market of 16%.


The company tax rate is 30%.


Required


3.1 Calculate the weighted average cost of capital using Gordon growth model to


calculate the cost of equity.

1
Expert's answer
2022-03-30T10:59:03-0400

"CARM=7+1.2(16-7)=17.8"


"r1=\frac{0.4(1+0.09)}{2.5}+0.09=0.2644"


"d2=2\times0.08=0.16"


"r2=\frac{0.16(1+0.09)}{1.8}+0.09=0.1869"


"capital=2 000 000\times1+1 500 000\times2+2 000 000 +10 000 000+1 000 000=18 000 000"


"WACC=\frac{2 000 000}{18 000 000}\times26.44+\frac{3 000 000}{18 000 000}\times18.69+\frac{10 000 000}{18 000 000}\times17.8+\frac{2 000 000}{18 000 000}\times10(1-0.3)+\frac{1 000 000}{18 000 000}\times15(1-0.3)=17.30"